OpenStudy (highschoolmom2010):

you see a rock climber on a cliff at a 32 degree angle. your eve level is 6 feet from the ground and you stand 1000 feet from the base of the cliff. what is the approx. height of the rock climber from the ground?

OpenStudy (highschoolmom2010):



OpenStudy (ankit042):

You know how to write sin for a triangle?

OpenStudy (highschoolmom2010):

yes

OpenStudy (highschoolmom2010):

\[\sin \frac{ opp }{ hyp }\]

OpenStudy (ankit042):

Yes try using this here we know sin30=1/2

OpenStudy (highschoolmom2010):

where did you get that from

OpenStudy (ankit042):

sin30 is a standard angle...It is good to know value of sin 0,30,45,60,90 otherwise you can always use calculator :P

OpenStudy (highschoolmom2010):

oh yea ok

OpenStudy (ankit042):

Now can you proceed from here?

OpenStudy (highschoolmom2010):

eh no not really

OpenStudy (ankit042):

where are you stuck?

OpenStudy (highschoolmom2010):

basically the begining this confused me bc i was given the adjacent side to the angle

OpenStudy (ankit042):

Sorry I couldn't understand you...can you explain more

OpenStudy (anonymous):

you use cos not sin if you are given adjecent. But better yet this problem requires tan :P

OpenStudy (anonymous):

tan because you are given the adjecent. And wanna know the oppossite side.

OpenStudy (anonymous):

tan(30)=opp/adj=x/1000

OpenStudy (highschoolmom2010):

so i need to use tan

OpenStudy (anonymous):

Yup :D it is the simplest trig ID that has both opp and adj in it.

OpenStudy (anonymous):

tan(30)=opp/adj=x/1000 tan(30)=x/1000 solve for x can you do that?

OpenStudy (highschoolmom2010):

\[1000*\tan 30=\frac{ x }{ 1000 }*1000\] \[1000( \tan 30)=x\] but the angle is 32 degrees

OpenStudy (anonymous):

1000*tan 32 =x is right so far :D what you do now is evaluate tan 32 in degree mode SOrry at typo

OpenStudy (anonymous):

i must have look at that other guy who say sin 30 lol.

OpenStudy (highschoolmom2010):

1000 tan (32) gave me a crazy answer

OpenStudy (anonymous):

1000*(tan 32) =x ok so we got this far all we do is use our calculator making sure we in degree mode not radian mode tan32=0.62486935190932750978051082794943665831087020284270496 so we jsut plug that in 1000*.6248=624.8 ft sounds believeable. ooh plus 6 feet for the guys height at eye level 624.8+6

OpenStudy (highschoolmom2010):

so rock climber is approx. 630.8' from the ground

OpenStudy (anonymous):

if that is one of your answer choices :D use the one that is closest to our answer. oh and i would go with that lol.

OpenStudy (ankit042):

Ohh sorry I made a mistake...was multitasking :(

OpenStudy (highschoolmom2010):

i have no answer choices

OpenStudy (highschoolmom2010):

@timo86m

OpenStudy (anonymous):

go with 630.8 then